OpenStudy (anonymous):

write 3 different expression that could be used to find the perimeter of the rectangle below

OpenStudy (solomonzelman):

This is how you find the perimeter of a rectangle. \(\huge\color{magenta}{ \rm L }\) \(\huge\color{darkviolet}{ \rm W }\) \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\Huge\frac{~~~~~~~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\) \(\huge\color{darkviolet}{ \rm W }\) \(\huge\color{magenta}{ \rm L }\) \(\huge\color{blue}{ \rm P_{rectangle} = \color{magenta}{ \rm 2L }+2\color{darkviolet}{ \rm W } }\)

OpenStudy (solomonzelman):

you can re-write it by factoring out of 2, or you can just expand the 2, like 2L is L+L, and do the same for W....

Nnesha (nnesha):

OMG!!!!!!!!!!!

Nnesha (nnesha):

@SolomonZelman you all know soooooooooooooooooooooooooooooooo manythings about latex and that rectangle

OpenStudy (solomonzelman):

it is just a trig of using the fractions of a same color as the font of the box, inside the box.